OpenStudy (anonymous):

Bob is wrapping gift boxes. Each gift box is a rectangular prism with a square base of side 9 inches. The difference between the heights of the two gift boxes is 6 inches. If the larger box has a surface area of 486 square inches, what is the difference in the surface area of the two boxes? Answer 54 square inches 108 square inches 144 square inches 216 square inches

OpenStudy (anonymous):



OpenStudy (anonymous):

 486= 2*(9*9)+4*(9*(x+6)) sfx x=3 Large box So now take difference 486-( 2*(9*9)+(4*(3*9))=216